two cars with speed of 90 kph and 72 kph travelling parallel .assuming each car to be 5m long find the time taken during the overtake and the total road distance used for the overtake
 

Grade:11

2 Answers

Saurabh Koranglekar
5 years ago
Dear student

90 kmph= 25mps

72kmph = 20 mps

initial difference is not given who is in lead is missing

please revert with complete question
Khimraj
5 years ago
I am assuming initial difference zero.
90 kmph= 25mps72kmph = 20 mps
Then relative distance = 5m
relative velocity = 25-20 = 5m/s
then time taken = 5/5 = 1sec
and distance travelled = 1*25 = 25 m